An organic compound A made up to 87.8% Carbon and 12.2% Hydrogen; reacts with cold dilute KMnO4 to yield D (C6H12O2). A takes up only one mole of hydrogen gas to form B when hydrogenated. A also decolourises aqueous Bromine to form C. (RAM of C = 12, H = 1.00, O = 16)

  • Determine the molecular formula of A. Hence give the structures of A, B, C & D
  • Compound C and D exist in several configurations. Draw the structures of any two configurations of either C or D.
1
Expert's answer
2021-03-10T05:58:51-0500

Moles of carbon = 87.8 / 12 = 7.32

Moles of Hydrogen = 12.2 / 1 = 12.2


Both are devided by 3.5 then


Moles of carbon = 7.32 / 3.5 = 2.02

Moles of hydrogen = 12.2 / 3.5 = 3.5


Moar ratio of carbon and hydrogen = 4 : 7



Emperical formula = C4 H7 .